OpenStudy (study.geek.123):

The equation of a circle is x^2+y^2+4x+8y=12. What is the center of the circle?

OpenStudy (hyuna301):

general eq. of the circle is x^2 + y^2 + 2gx + 2fy + c = 0 and center of the circle is ( -g , -f ) so, center = ( -2, -4)

OpenStudy (mww):

You will need to convert the equation to the centre, radius form. x^2+y^2+4x+8y=12 You must gather the x's and the y's separately and complete the square for each x^2+y^2+4x+8y=12 x^2+4x + y^2+8y=12 (move x's together, y's together) Now we need to complete the square for both x and y. a^2 + 2ab + b^2 = (a + b)^2 Remember to do this you find half of the coefficient of x or y, then square that and add it at the end. x^2+4x+4 + y^2 + 8y + 16 = 12 + 4 + 16 (remember to balance both sides) (x+2)^2 + (y + 4)^2 = 32 (rewrite as perfect square) So the centre is (-2,-4) and the radius is sqrt(32) = 4 sqrt(2)